OpenStudy (anonymous):

[7.08] Find the quotient of quantity of 24 times x to the 2nd power times y plus 12 times x times y to the 3rd power minus 18 times x times y all over 6 times x times y. Select one: a. 6y2 − 18 b. 4x + 2y2 − 3 c. 12x2 − 18x + 3 d. x3y2 + 2x2y4 − 4x2y2

OpenStudy (johnweldon1993):

\[\large \frac{24x^2y + 12xy^3 - 18xy}{6xy} = \frac{24x^2y}{6xy} + \frac{12xy^3}{6xy} - \frac{18xy}{6xy} = \space ?\]

OpenStudy (anonymous):

I'm not sure, d?

OpenStudy (johnweldon1993):

Lets do it by parts... \[\large \frac{24x^2y}{6xy} = \frac{24}{6} \times \frac{x^2}{x} \times \frac{y}{y} = 4 \times x \times 1 = 4x\] Thats the first part...can you do the rest?